Write down wavefunction for state n=1
l=0 m=0
1
Expert's answer
2020-11-02T11:08:51-0500

The wave function for the state e n=1 l=0 m=0 is

"\psi_{100}=\frac{1}{\sqrt \pi}(\frac{1}{a_0})^{2/3}e^{-\frac{r}{a_0}}"

Here "a_0" is the Bohr radius.